OpenStudy (anonymous):

Solve: (1/4)(12 − 4t) = 27 -24 24 -6 6

OpenStudy (nali):

first multiply 1/4 by 12 and then by -4t

OpenStudy (nali):

and then just solve for t

OpenStudy (nali):

need any additional help

OpenStudy (anonymous):

im still not getting the answer

OpenStudy (anonymous):

\(\large \frac{1}{4}(12-4t)=27\rightarrow(12-4t)=(27)(4)\rightarrow -4t=(27)(4)-12\) \(\large t = \frac{(27)(4)-12}{-4}\)